Write each equation in ax+by=c form.
1. y=5-3x
for this one, I got y-3x=5 , is that right?
2. 2/3x+1/4

Answer by ankor@dixie-net.com(22740)

1. y = 5 - 3x
add 3x to both sides and you have
3x + y = 5, which is the form that you want
:
Assume it's
2. y = (2/3)x + 1/4
we want integer coefficients here
multiply by 12
12y = 8x + 3
subtract 8x from both sides
-8x + 12y = 3